OpenStudy (anonymous):

The tax on a luxury item in Kansas is computed based upon 2% of the first $1000, 4% of the next $2000 and 5% thereafter. What would the tax be on a luxury item valued at $30,000? $200 $800 $1350 $1450 $5700

OpenStudy (anonymous):

30000 = 1000 + 2000 + 27000 TotalTax= 0.02*1000+0.04*2000+0.05*27000 TotalTax= $1450